2.1. Solve 2xy + 6x + (x^2 - 4)y'=0


1
Expert's answer
2022-06-27T06:38:19-0400
"(x^2 - 4)y'+2xy=-6x"

"y=uv\\
y'=u'v+uv'"

"(x^2-4)(u'v+uv')+2xuv=-6x"

"u((x^2-4)v'+2xv)+(x^2-4)vu'=-6x"

Let

"(x^2-4)v'+2xv=0"

Then

"\frac{dv}{v}=-\frac{2x}{x^2-4}dx"

"\ln v=-\ln (x^2-4)"

"v=\frac{1}{x^2-4}"

We get an equation

"(x^2-4)\frac{1}{x^2-4}u'=-6x"

"u'=-6x"

"u=-3x^2+C"

Finally

"y=uv=\frac{-3x^2+C}{x^2-4}"